# Traffic Response Functions: Patterns, Propagation and Congestion

### v1

@@ Abstract
Using empirical data gathered on motorways in Germany, we follow a new approach by further exploring response functions as a possible tool to study traffic dynamics in motorway networks. We uncover the basic characteristics of responses of flow and density to given signals and the capability of responses to capture the correlation between these fundamental observables. Furthermore, we uncover the potential use of responses to characterize traffic patterns. We are able to demonstrate the differentiation of congestion patterns and the determination of the propagation velocity of moving congestion.
@@ 1 Introduction
Vehicular traffic on a multi-lane motorway is a highly complex and dynamic process. Nevertheless, millions of traffic participants experience this complexity on daily basis due to a rather simple fact that seems to be inevitable, namely the fact of traffic congestion. Although it may seem very simple to the individual driver, congestion is a phenomenon emerging in a collective driven many-body system which unfolds its complex dynamics in different scales of space and time. Over the last six decades a large variety of theories and models has been developed to understand and describe the complex, non-linear behavior of traffic systems, e.g. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] and others. Due to the fact that vehicular traffic is a key element of modern society’s mobility and transportation it is a desirable task to continuously study traffic networks. The development of alternative research approaches and methods helps to understand these systems and their dynamics much better and find new tools for effective traffic planning, management and control.
Recently, we successfully conducted several studies along these lines [16, 17, 18, 19, 20], focusing on the application of concepts and approaches for studying financial markets on traffic networks. By extensively studying traffic data with correlation matrices, we identified dominant and subdominant collectivities in traffic, revealing new characteristics of traffic networks. The most recent work in [20] uncovers the potential application of response functions as a tool for studying traffic dynamics. The basic idea behind the concept of response functions is being able to study a systems reaction to specific events, revealing their influence on and the sensitivity of the response variable. The applications of response functions in financial markets have successfully revealed the influence of different time scales on non-Markovian effects [21, 22, 23, 24, 25, 26]. In case of vehicular traffic, congestion on motorways is certainly one of the most interesting and significant events to which responses of traffic observables can be studied. By choosing the velocity as a response variable, response functions are able to capture the presence of congestion, its propagation along motorway segments and therefore spatiotemporal features of present traffic patterns [20]. These findings lay an important foundation for the study of traffic networks with the concept of response functions. However, the attempt of fully characterizing traffic patterns is a difficult but crucial task and information on other fundamental observables, such as density and flow, are desirable.
@@
With this contribution we aim to deepen the understanding of the possibilities of response functions as a tool for the study of traffic dynamics. For this sake, we extend the previous approach in [20] by also incorporating flow and density responses for spatially subsequent motorway sections. By comparing all three types of responses we aim to uncover the basic characteristics of flow and density responses. Furthermore, we define multiple congestion indicators according to different velocity ranges. The goal is to identify different phases or patterns of congestion. In the course of this we investigate the possibility of deriving the propagation velocity of congestion with the help of responses.
@@
This paper is organized as follows. Sec. 2 presents the studied data set and introduces response functions and associated indicators for congestion. Sec. 3 introduces the locations we use to conduct the study. The results are presented in Sec. 4. Lastly, Sec. 5 briefly summarizes and concludes our results.
@@ 2 Data description and methods
For our study we use traffic data gathered by inductive loop detectors on motorways in North Rhine-Westphalia (NRW), Germany. In the following we describe the used data in Sec. 2.1 and present the methods conducted in our study in Sec. 2.2.
@@ 2.1 Data
Inductive loop detectors on motorways gather information about the traffic situation by recording essential observables on each lane, such as traffic flow rate $q$ and average velocity $v$ [27]. At each section the measurements are conducted over time intervals of one minute and the aggregated data are given in units of vehicles/h and km/h, respectively. In addition, the detectors distinguish cars and trucks. Our study focuses on workdays, therefore we exclude weekends and public holidays, leaving a total number of $N=243$ workdays from our data set recorded from December 2016 to December 2017.
For the purpose of this study we combine the data of all lanes at each section into one effective lane as follows. For each lane $l$ at each section $s$ we determine a traffic density
$\rho_{sl}^{\mathrm{(m)}}(t)=\frac{q_{sl}^{\mathrm{(m)}}(t)}{v_{sl}^{\mathrm{(m)}}(t)},$ (1)
with $\mathrm{m}=\text{car},\text{tr}$, for cars and trucks separately. The sum of the respective quantities of cars and trucks yield the total traffic flow and density
@@
for each lane. The total, effective flow $q_{s}(t)$ and density $\rho_{s}(t)$ for each section result from the summation over all lanes. Finally, the corresponding effective velocity $v_{s}(t)$ at each section is given by
@@
Even though using the hydrodynamic relation in Eq. (1) is a common procedure for the determination of traffic densities from loop detector data, we like to point out, for the sake of completion, that it has to be considered as an approximation. Here we determine a spatial quantity from a temporal quantity, which can result in systematic errors [27]. By using the time mean speed for calculation with Eq. (1), instead of the spatial mean speed, the true density is usually underestimated. However, in most cases the approximation is considered sufficient because the main characteristics of macroscopic traffic effects are preserved.
@@ 2.2 Response functions and congestion indicators
where $\tilde{R}_{ij}^{(x)}(\tau)$ yields the response, i.e. the average change, of a traffic observable $x$ at section $i$ at time $t+\tau$ conditioned on congestion indicated by $\varepsilon_{j}(t)$ at section $j$ at a previous time $t$. With regard to [20], we take a first step to extend the study by not only considering the responses of velocities $x=v$, but also determining and discussing the responses of flow $x=q$ and density $x=\rho$. In a second step, we take into account the complex nature of freeway traffic by varying the indicator $\varepsilon_{j}(t)$. Alternatively one could also define response functions in which the unconnected part $\left\langle\Delta x_{i}(t,\tau)\right\rangle_{t}\left\langle\varepsilon_{j}(t)\right\rangle_{t}$ is subtracted. Here we find it useful to use the definition in Eq. (5).
On the one hand, in a relatively simple approach congestion can be described as a (density) shock wave propagating up-stream through free flow traffic phases [1, 3]. On the other hand, more recent traffic theory approaches show the existence of different phases (or states) of congestion and their transitions, whose description is far more complicated, e.g. three-phase traffic theory [14, 15]. In light of the research in [20], we have an idea of the properties that response functions may capture in relation to congestion phases. We expect to find the capability of responses to reflect the correlations between the fundamental traffic observables velocity, flow and density during phases of congestion propagation as depicted in Fig. 1.
@@
We try to deepen the study on traffic dynamics by using response functions with four different indicator functions
@@
More than one indicator function help to capture different phases of congestion occurrence such as stop-and-go traffic, wide moving jams, and others. In this way, we also show an example for the flexible applications of response functions according to different research purposes. While $\varepsilon^{(0)}(t)$ indicates heavily congested sections with low velocity range in Eq. (6), $\varepsilon^{(1)}(t)$ and $\varepsilon^{(2)}(t)$ aim to indicate situations like stop-and-go traffic or phases between two congestion waves. Incorporating the velocity ranges of the previous indicators, $\varepsilon^{(3)}(t)$ represents a more generous threshold for the indication of congestion. It is mainly introduced for the sake of comparison.
Our work focuses on rush hours to study the traffic responses to congestion, more specific on either morning rush hours (mr) from 06:00 to 11:00 or afternoon rush hours (ar) from 14:00 to 19:00, depending on the individual cases to be examined. The time lag in Eq. (5) is restricted to $\tau=300\ \text{min}$ covering the whole time interval of the rush hours. To avoid capturing responses to random events, e.g. accidents or daily construction sites, in this study response functions are averaged over the given work days. We generally proceed as follows: For each case examined, we start by choosing an indicator section $j$ and convert all given time series of velocities $v_{j}(t)$ into four times series of indicators $\varepsilon_{j}^{(y)}(t)$, $y=0,1,2,3$ according to Eq. (6). For each day and each pair of sections $ij$ the corresponding response functions $\tilde{R}_{ij}^{(x)}(\tau)$ are calculated for all three observables velocity, flow and density from section $i$ for all four types of indicators from section $j$. Averaging the calculated responses over all given days results in the averaged response function $R_{ij}^{(x)}(\tau)$, $x=v,q,\rho$, respectively, for each given indicator.
@@ 3 Studied locations
Motorway interchanges are important junctions in traffic networks and usually provide more than one function. They establish a connection between two (or more) motorways, they allow traffic participants to change the course of direction and, in some cases, they function as an entry point into the motorway network. Unfortunately, as useful as they are, interchanges also have downsides during heavily trafficked periods. Due to their structure, combined with the fact that on- and off-ramps act as bottlenecks [14, 15], motorway interchanges are very likely to be an origin of traffic congestion during rush hours. This is especially the case for motorway interchanges connecting regularly high trafficked motorways. Apart from inducing traffic congestion, motorway interchanges are capable of spreading congestion into other parts of the motorway network they connect with. As a result, congestion may propagate through on- and off-ramps onto connected motorways or urban traffic networks, which subsequently leads to a deterioration of the present traffic situation.
@@
Figure 2(a) shows the motorway interchange Breitscheid, which connects three motorways, namely A52, A3 and A524. Motorway A3 is a longer ranged motorway directly connecting big cities like Cologne (in south direction), Duisburg and Oberhausen (in north direction). Motorway A52 is a shorter ranged motorway which establishes a direct connection between the cities Düsseldorf (in south-west direction) and Essen (in north-east direction). As shown in Fig. 3(a), the segment of A52 in north-east direction, going from Düsseldorf to Essen, is heavily trafficked during rush-hours and congestion is present on daily basis. In addition, the traffic situation is complicated because traffic participants are able to leave and enter the motorway network before passing section 6 and section 2, respectively. Section 6 is a special case due to the general structure of the interchange Breitscheid. It is located on a separated lane which routes the traffic flow from connected motorways, merging it with the traffic on the main lanes of A52 between section 2 and section 1 (see supplementary map in Fig. A1 in Appendix A.1). For the conducted investigations section 1 will be the indicator section.
Motorway interchange Herne, shown in Fig. 2(b), connects motorway A42 and A43. While the former provides a connection of large cities of the Ruhrgebiet (Ruhr area), such as Gelsenkirchen, Bottrop, Oberhausen and Duisburg (from east to west), the latter connects the city of Herne with Wuppertal, Bochum, Recklinghausen and Münster (from south to north). The overall amount of traffic within interchange Herne is not quite as high as in the case of interchange Breitscheid, but Fig. 3(b) clearly demonstrates the occurrence of congestion on regular bases during rush-hours. Depending on the study case, sections 11 and 9 will be the indicator sections for our investigation.
The third location is the segment of motorway A3 between interchanges Breitscheid and Kaiserberg, see Fig. 2(c). This part of the A3 is heavily trafficked in north direction during afternoon rush-hours on daily basis. Within interchange Kaiserberg, near section 14 and (spatially) right next to an off-ramp, the number of main lanes reduces from three to two. This reduction acts as an additional bottleneck, reducing the capacity of the motorway and inducing heavy congestion on regular basis. Figure 3(c) demonstrates how severe the congestion is on average and how far the traffic breakdown propagates into the south. Due to the traffic situation within interchange Kaiserberg, we choose section 14 as the indicator section for our investigation.
@@ 4.1 A comparison of responses of velocity, flow and density
A convenient starting point for our analysis is to briefly relate this study to previous results from [20], in which interchange Breitscheid was already a location of investigation. Figure 4(a) shows velocity, flow and density responses of spatially subsequent sections to congestion at section 1 for all used indicators during afternoon rush hours. Despite a different velocity range of indicator $\varepsilon^{(0)}(t)$ for heavy congestion (up to 20 km/h instead of 10 km/h used in [20]), one can easily identify similar behavior of velocity responses by comparison. The responses of neighboring sections 2 to 5 show negative values, forming dipping-downs or local minima, for smaller time lags and increase to positive values, continuously growing for greater $\tau$. The occurrence of response valleys, as well as their amplitude and width, depend on the distance between indicator section $j$ and the responding section $i$. With increasing distance the valleys occur at greater time lag with shrinking width and amplitude. This response behavior captures and reflects the velocity breakdown and therefore the upstream propagation of congested traffic. By comparison, we find the behavior of velocity responses for $\varepsilon^{(0)}(t)$ at interchange Breitscheid is the same as the ones at interchange Herne in Fig. 4(b) and on motorway A3 in Fig. 4(c). This demonstrates that the velocity responses for $\varepsilon^{(0)}(t)$ are independent of the studied location.
@@
Before we discuss the response for other indicators, we take a first look at flow and density responses for indicator $\varepsilon^{(0)}(t)$ in Fig. 4. Even though the flow has a very fluctuating nature on shorter time scales (a couple of minutes), the average responses show very clear courses, including sharp local minima which occur at larger time lags with increasing distance to the indicator section. This is the case for all three studied locations in Fig. 4. The amplitudes of the negative responses of the nearest neighboring sections to the indicator section at Breitscheid ($i=2$) and motorway A3 ($i=15$) in Fig. 4(a) and (c) corroborate the severity of the occurring congestion at these locations. Values between -1000 veh/h and -1200 veh/h clearly demonstrate the heavy traffic breakdown already indicated by the average data in Fig. 3. A comparison of velocity and flow responses with the density responses in Fig. 4 reveals fulfilled expectations regarding their courses. Due to the fact that traffic breakdowns induce a local increase in density, the respective density responses actually show local maxima which occur at larger time lags $\tau$ for increasing distance to the indicator section. Also, in case of density response the scaling of the amplitudes reflect the severeness of the traffic jam at Breitscheid and on A3 in Fig. 4(a) and (c). With values over 20 veh/km for the nearest responding section the local increase in density amounts to a fourth of the average value shown in Fig. 3. Overall, we find that the local minima and maxima of all response types occur at the same time lag $\tau$ for the individual responding sections. Therefore, the responses of all three observables for $\varepsilon^{(0)}(t)$ clearly capture the propagation of congestion and the correlations between velocity, flow and density as depicted in Fig. 1. This finding is a fundamental basis for response functions as a tool to study traffic dynamics.
@@
Taking a closer look at the responses for the remaining congestion indicator, one also finds the above described correlations satisfied for all three observables with indicator $\varepsilon^{(2)}(t)$ being a special case. Before discussing the latter, we begin with the more generous velocity range of indicator $\varepsilon^{(3)}(t)$, which demonstrates the importance of choosing smaller velocity ranges to indicate congestion when studying traffic dynamics with response functions. While the minima and maxima of responses to $\varepsilon^{(3)}(t)$ are visible in case of interchange Breitscheid and motorway A3 in Fig. 4(a) and (c), the response at interchange Herne does not exhibit any clear maxima or minima in Fig. 4(b). Thus, it would not be possible to obtain any information about the traffic situation at interchange Herne. In addition, the amplitudes of the local minima and maxima of the responses to $\varepsilon^{(3)}(t)$ show either much smaller values, or values in range of indicator $\varepsilon^{(2)}(t)$ in case of motorway A3 and interchange Breitscheid. Both findings demonstrate that one is able to extract more detailed information about the system by applying multiple indicators. If we restrict a comparison of the remaining three indicators to the case of interchange Breitscheid, we find the amplitudes of the minima (maxima) shrinking (increasing) with increasing values of the velocity ranges. In terms of traffic theory one may expect to find this behavior. During formation of a dense congested phase, i.e. small distances between vehicles moving with small average velocity (e.g. $v<20$ km/h ), an abrupt deceleration to small velocities induces a strong drop of flow rate and increase of local density, respectively. In contrast, less dense congestion allows vehicles to move with higher velocities, resulting in smaller drops of flow rate and the decrease of local density. Under the assumption of vehicle count preservation, the hydrodynamic relation in Eq. (1) incorporates this.
@@
The response to indicator $\varepsilon^{(2)}(t)$ shows a behavior opposite to the one of other responses at smaller $\tau$ in case of interchange Herne and motorway A3, see Fig. 4(b) and (c). Instead of finding local minima for velocity and flow response, and local maxima for density response, respectively, the responses to indicator $\varepsilon^{(2)}(t)$ exhibit local maxima in case of velocity and flow, and local minima in case of density. This opposite behavior is a result of positive increments $\Delta x_{i}(t,\tau)$ in Eq. (5) and represents vehicles accelerating during the congestion phase, increasing the flow and lowering the local density. Consequently, the correlations between the three observables remain, but here they express themselves in the opposite case. However, the results depicted in Fig. 4 indicate a difference between the phases of congestion occurring at interchange Breitscheid and the ones occurring at interchange Herne and on motorway A3. Besides the acceleration phase captured by $\varepsilon^{(2)}(t)$, we find a similar case in the flow response for $\varepsilon^{(0)}(t)$ at interchange Breitscheid in Fig. 4(c). A few minutes after the minimum, the response to $\varepsilon^{(0)}(t)$ transitions into a local maximum indicating vehicle acceleration out of the congested traffic. Unfortunately, this only manifests in the flow response and not in velocity or density responses to $\varepsilon^{(0)}(t)$. Although no distinct local maxima occur, there is an increasing course in both responses. Overall, response functions are capable of capturing acceleration phases during congestion, which is an important finding. It can be considered as another foundation for their capability as a tool to study traffic dynamics, especially due to the fact that one is able to identify differences in congestion phases. In addition, it further demonstrates the importance of choosing several indicators in terms of velocity ranges.

### v2

@@ Abstract
We further pursue a recently tested approach to the study of transport systems. Using empirical data collected on German motorways, we investigate the use of response functions as a tool to study traffic dynamics in motorway networks. We use a response function that measures the change in traffic observed at a given location over time conditioned on congestion at another location. By introducing several velocity-dependent congestion indicators, we are able to work out the basic characteristics of the flow and density responses to congestion. Our results demonstrate that these responses capture the propagation of congestion, the acceleration phases of traffic participants during congestion, and the correlation between the three fundamental traffic observables: flow, density, and velocity. Furthermore, a direct comparison with the underlying data reveals the potential for characterizing traffic patterns with response functions. A relationship between the shape of the response functions and the traffic oscillation during a traffic jam can be demonstrated. Moreover, our findings allow us to determine the propagation velocity of congestion from responses. Importantly, in the course of our analysis, we also identify a scaling property that results in the collapse of the response functions onto a single curve.
@@ 1 Introduction
Vehicular traffic on a multi-lane motorway is a highly complex and dynamic process. Millions of drivers experience this complexity on a daily basis due to a rather simple and seemingly unavoidable fact: traffic congestion. Congestion is a phenomenon arising in a collectively driven many-body system that unfolds its complex dynamics in different scales of space and time. Over the past six decades, a variety of theories and models have been developed to comprehend and describe the complex, non-linear behavior of traffic systems. These include fluid-dynamical theories [1, 3], gas-kinematic models [2, 5], car-following theories [4, 6, 11], cellular automata models [7, 8, 9, 10], as well as the three-phase traffic theory [12, 13, 14], and others. In view of the central role played by vehicular traffic in the mobility and transportation of modern societies, alternative research methodologies may facilitate a more comprehensive understanding of these systems and their dynamics, thereby enabling the development of new tools for effective traffic planning, management and control.
We have recently conducted a number of studies along these lines, focusing on the application of concepts and approaches used to study financial markets in the context of traffic networks. We conducted a comprehensive analysis of traffic data to investigate temporal [15] and spatial [17] correlations, leading to the identification of dominant and subdominant collectivities [16, 18] and new characteristics of traffic networks. The most recent work [19] explores the potential of response functions as a tool for studying traffic dynamics. The fundamental premise of the concept of response functions is the ability to examine a system’s reaction to specific events, thereby revealing the influence of these events on the response variable and the sensitivity of the system to such influences.
In the field of finance, response functions have been successfully applied to analyse the reaction of a system to specific events on different time scales, highlighting the complexity of market behaviors. Previous trades and correlated sector responses drive present and future price dynamics. In this context, response functions have enabled significant advances in our understanding of cross-impact mechanisms [21, 23, 24], liquidity effects [23, 25], sectoral dependencies [22], and the non-Markovian nature of price fluctuations [20] in financial markets. In the context of transportation networks, congestion is one of the most significant events to study the responses of traffic observables. By selecting velocity as the response variable, response functions can identify the presence of congestion, its propagation along motorway segments, and therefore spatiotemporal characteristics of traffic patterns [19]. These findings establish a crucial foundation for studying traffic networks with the concept of response functions. Full characterization of traffic patterns is a challenging task, and information on other fundamental observables, such as density and flow, is valuable.
@@
The aim of this contribution is to further establish response functions as a research instrument for the study of traffic dynamics. To this end, we extend the previous approach presented by Wang et al. [19] by also incorporating flow and density responses for spatially subsequent motorway sections. By comparing all three types of responses, our objective is to identify the fundamental characteristics of flow and density responses. Additionally, we define multiple congestion indicators according to different velocity ranges. The aim is to identify different phases or patterns of congestion. In this process, we also investigate the potential for deriving the propagation velocity of congestion using responses.
@@
The paper is organized as follows. Sec. 2 presents the data set studied and introduces response functions and related indicators of congestion. Sec. 3 introduces the locations used for the study. The results are presented in Sec. 4. Sec. 5 briefly summarizes and concludes our results.
@@ 2 Data description and methods
For our study we use traffic data collected by inductive loop detectors on motorways in North Rhine-Westphalia (NRW), Germany. In Sec. 2.1 we describe the data used for the study and present the methods applied in Sec. 2.1.
@@ 2.1 Data
Inductive loop detectors deployed on motorways are designed to gather information about the traffic situation by recording essential observables on each lane [26]. These observables are the traffic flow rate $q$ and the average velocity $v$. The measurements are conducted at each section over time intervals of one minute, and the aggregated data are presented in units of vehicles/h and km/h, respectively. Furthermore, the detectors are capable of differentiating between cars and trucks. Our study is limited to workdays, (weekends and public holidays are excluded) resulting in a total of $N=243$ workdays from December 2016 to December 2017 included in our data set.
In this study the data from all lanes at each section are aggregated as a single effective lane. For each lane $l$ at each section $s$ a traffic density
$\rho_{sl}^{\text{(m)}}(t)=\frac{q_{sl}^{\text{(m)}}(t)}{v_{sl}^{\text{(m)}}(t)},$ (1)
[…]
@@
for each lane. The total effective flow $q_{s}(t)$ and density $\rho_{s}(t)$ for each section are obtained by summing over all lanes. The corresponding effective velocity $v_{s}(t)$ at each section is then given by
@@
Although the use of the hydrodynamic relation in Eq. (1) is a standard procedure for determining traffic densities from loop detector data, it should be noted that this method is an approximation. Deriving a spatial quantity from a temporal quantity can potentially introduce systematic errors [26]. By employing the time averaged velocity in the calculation with Eq. (1), rather than the spatially averaged velocity, the actual density is typically slightly underestimated. Nevertheless, in the majority of cases, this approximation is deemed sufficient since the primary characteristics of macroscopic traffic effects are preserved.
@@ 2.2 Response functions and congestion indicators
which yields the response, or the average change, of a traffic observable $x$ at section $i$ at time $t+\tau$ conditioned on congestion indicated by $\varepsilon_{j}(t)$ at section $j$ at a previous time $t$. In regard to the study by Wang et al. [19], we take the initial step of extending the analysis to encompass not only the responses of velocities ($x=v$), but also those of flow ($x=q$) and density ($x=\rho$). In a second step, the complex nature of freeway traffic is taken into account by varying the indicator $\varepsilon_{j}(t)$. An alternative approach would be to define response functions in which the unconnected part $\left\langle\Delta x_{i}(t,\tau)\right\rangle_{t}\left\langle\varepsilon_{j}(t)\right\rangle_{t}$ is subtracted. Here we find it useful to use the definition in (5).
On the one hand, in a relatively simple approach congestion can be described as a (density) shock wave propagating upstream through free flow traffic phases [1, 3]. In contrast, more recent traffic theory approaches demonstrate the existence of diverse phases (or states) of congestion and their transitions, which are more complex to describe, e.g. three-phase traffic theory [13, 14]. Considering the results of Wang et al. [19], we have developed an understanding of the characteristics that response functions may encompass in relation to congestion phases. Our analysis will reveal the capacity of responses to reflect the correlations between the fundamental traffic observables, namely velocity, flow, and density, during phases of congestion propagation, as illustrated in Fig. 1.
@@
As compared to the response functions in Ref. [19], we here refine the analysis by conditioning on velocity intervals. To this end, we analyse responses that incorporate the four different indicator functions
@@
Fine-tuned indicator functions enable the capture of different phases of congestion occurrence, such as stop-and-go traffic, wide moving jams, and others. In Eq. (6), $\varepsilon^{(0)}(t)$ indicates heavily congested roadways with low velocity range. Indicators $\varepsilon^{(1)}(t)$ and $\varepsilon^{(2)}(t)$ are designed to indicate situations such as stop-and-go traffic or phases between two congestion waves. The introduction of the velocity ranges of the previous indicators results in $\varepsilon^{(3)}(t)$, which represents a more inclusive threshold for the indication of congestion for comparison.
We focus on the analysis of traffic responses to congestion during rush hours. Specifically, we examine either morning rush hours (mr) from 06:00 to 11:00 or afternoon rush hours (ar) from 14:00 to 19:00, depending on the individual cases to be examined. The time lag in Eq. (5) is constrained to a maximum of $\tau=300\ \text{min}$, covering the entirety of the rush hour interval. To avoid random events like accidents or construction sites affecting the results, we average response functions over the specified work days in this study. We proceed as follows: For each case under examination, an indicator section $j$ is selected, and all given time series of velocities $v_{j}(t)$ are transformed into four time series of indicators $\varepsilon_{j}^{(y)}(t)$, with $y=0,1,2,3$, according to Eq. (6). For each day and each pair of sections $ij$ the corresponding response functions $\tilde{R}_{ij}^{(x)}(\tau)$ are calculated for all observables (velocity, flow and density) for all four types of indicators. In each case, the calculated responses are averaged over all workdays, yielding the averaged response function $R_{ij}^{(x)}(\tau)$, with $x=v,q,\rho$ respectively, for the given indicator.
@@ 3 Studied locations
Interchanges are important nodes in transportation networks and typically serve more than one function. They provide a connection between two (or more) motorways, allow travelers to change direction, and in some cases serve as an entry point to the motorways network. They also have drawbacks during heavy traffic periods. Due to their structure, combined with the fact that on- and off-ramps act as bottlenecks [13, 14], motorway interchanges are very likely to be a source of traffic congestion during rush hours. This is particularly true for interchanges that connect highly trafficked motorways. In addition to inducing congestion, interchanges have the potential to spread congestion to other parts of the motorway network. Congestion can spread through on-ramps and off-ramps to connected motorways or urban traffic networks, resulting in a deterioration of the current traffic situation.
@@
Figure 2(a) shows the Breitscheid interchange, which connects three motorways, namely A52, A3 and A524. The A3 is a longer-range motorway that connects large cities such as Cologne (in the south), Duisburg and Oberhausen (in the north). Motorway A52 is a shorter-range motorway that connects the cities of Düsseldorf (in southwest direction) and Essen (in northeast direction). As shown in Fig. 3(a), the segment of the A52 in northeast direction, going from Düsseldorf to Essen, is heavily trafficked during rush hours and congestion is present on a daily basis. The traffic situation is complicated by the fact that traffic participants can exit and enter the motorway network before passing section 6 and section 2, respectively. Section 6 is a special case due to the general structure of the Breitscheid interchange. It is located on a separate lane which routes the traffic flow from the connected motorways and merges it with the traffic on the main lanes of A52 between section 2 and section 1 (see supplementary map in Fig. A1 in Appendix A.1). Section 1 will be the indicator section for the investigations performed.
The Herne interchange, shown in Fig. 2(b), connects motorway A42 and A43. While the former connects large cities of the Ruhrgebiet (Ruhr area), such as Gelsenkirchen, Bottrop, Oberhausen and Duisburg (from east to west), the latter connects the city of Herne with Wuppertal, Bochum, Recklinghausen and Münster (from south to north). The total amount of traffic within the Herne intergange is not quite as high as in the case of the Breitscheid interchange, but Fig. 3(b) clearly demonstrates the occurrence of congestion on a regular bases during rush hours. Depending on the case under study, sections 11 and 9 will be the indicator sections for our investigation.
The third location is the segment of motorway A3 between the Breitscheid and Kaiserberg interchanges, see Fig. 2(c). This part of the A3 is heavily trafficked northbound during afternoon rush hours on a daily basis. At the Kaiserberg interchange , near section 14 and (spatially) right next to an off-ramp, the number of main lanes is reduced from three to two. This reduction acts as an additional bottleneck, reducing the capacity of the motorway and causing heavy congestion on a regular basis. Figure 3(c) demonstrates how severe the congestion is on average and how far the traffic breakdown propagates to the south. Due to the traffic situation within the Kaiserberg interchange, we choose section 14 as the indicator section for our investigation.
@@ 4.1 A comparison of responses of velocity, flow and density
It is useful to briefly relate our study to previous results from Wang et al. [19], in which interchange Breitscheid was already investigated. Figure 4(a) depicts the velocity, flow, and density responses of spatially subsequent sections to congestion at section 1 for all employed indicators during afternoon rush hours. Despite a different velocity range for indicator $\varepsilon^{(0)}(t)$ in the case of heavy congestion (up to 20 km/h instead of 10 km/h used in Ref. [19]), it is possible to identify similar behavior of velocity responses by comparison. The responses of neighboring sections 2 to 5 exhibit negative values, manifested as local minima or dips, for shorter time lags and rise to positive values, displaying continuous growth for longer time lags $\tau$. The occurrence, amplitude, and width of response minimas depend on the distance between indicator section $j$ and the responding section $i$. As the distance increases, the minima occur at longer time lags with shrinking width and amplitude. This response behavior captures and reflects the velocity breakdown and, consequently, the upstream propagation of congestion. In comparison, the velocity responses for $\varepsilon^{(0)}(t)$ at interchange Breitscheid are found to be identical to those observed at interchange Herne in Fig. 4(b) and on the A3 in Fig. 4(c). This evidence suggests that the velocity responses for $\varepsilon^{(0)}(t)$ are independent of the specific location under study. We view this as an important identification of universalities in responses.
@@
Prior to examining the responses of other indicators, we will initially investigate the flow and density responses to $\varepsilon^{(0)}(t)$ in Fig. 4. Despite the inherently fluctuating nature of the flow on shorter time scales (a couple of minutes), the average responses exhibit a clear trend, including the occurrence of sharp local minima at larger time lags with increasing distance to the indicator section. This is observed in all three locations studied in Fig. 4. The amplitudes of the negative responses of the nearest neighboring sections to the indicator section at Breitscheid ($i=2$) and motorway A3 ($i=15$) in Fig. 4(a) and (c) corroborate the severity of the congestion occurring at these locations. Values between -1000 veh/h and -1200 veh/h clearly demonstrate the heavy traffic breakdown indicated by the average data in Fig. 3. A comparison of velocity and flow responses with the density responses in Fig. 4 reveals that the expected courses have been met. As traffic congestion causes an increase in density at the local level, the respective density responses exhibit local maxima, occurring at larger time lags $\tau$ for increasing distances from the indicator section. Furthermore, the scaling of the amplitudes in the density response reflects the severity of the traffic congestion at Breitscheid and on the A3, as illustrated in Fig. 4(a) and (c). With values exceeding 20 veh/km for the nearest responding section, the local increase in density represents approximately a quarter of the average value shown in Fig. 3. In summary, the local minima and maxima of all response types occur at the same time lag $\tau$ for the individual responding sections. Consequently, the responses of all three observables for $\varepsilon^{(0)}(t)$ capture the propagation of congestion and the correlations between velocity, flow, and density, as illustrated in Fig. 1.
@@
A more detailed examination of the responses for the remaining congestion indicator reveals that the aforementioned correlations are also present in all three observables, with indicator $\varepsilon^{(2)}(t)$ representing a special case. Before discussing the latter, we will begin with the discussion of indicator $\varepsilon^{(3)}(t)$, which demonstrates the importance of choosing smaller velocity ranges to indicate congestion when studying traffic dynamics with response functions. While the minima and maxima of responses to $\varepsilon^{(3)}(t)$ are visible in case of interchange Breitscheid and motorway A3 in Fig. 4(a) and (c), the response at interchange Herne does not exhibit any clear maxima or minima in Fig. 4(b). Thus, it would not be possible to obtain any information about the traffic situation at interchange Herne. In addition, the amplitudes of the local minima and maxima of the responses to $\varepsilon^{(3)}(t)$ show either much smaller values, or values in range of indicator $\varepsilon^{(2)}(t)$ in case of motorway A3 and interchange Breitscheid. These findings demonstarte that the application of multiple indicators enables the extraction of more detailed information about the system. When the comparison of the remaining three indicators is restricted to the case of interchange Breitscheid, it can be observed that the amplitudes of the minima (maxima) decrease (increase) with increasing values of the velocity ranges. In terms of traffic flow theory, this behavior is expected. During the formation of a dense congestion phase, i.e. small distances between vehicles moving at low average velocity (e.g. $v<20$ km/h ), an abrupt deceleration to low velocities induces a large drop in flow rate and an increase in local density. In contrast, less dense congestion allows vehicles to move at higher velocities, resulting in a smaller drop of flow rate and a decrease in local density. Under the assumption of vehicle number preservation, the hydrodynamic relation in Eq. (1) takes this into account.
@@
The response to indicator $\varepsilon^{(2)}(t)$ contrasts with the other responses at smaller $\tau$ in the case of interchange Herne and motorway A3, as illustrated in Fig. 4(b) and (c). Opposed to the expected behavior, the responses to indicator $\varepsilon^{(2)}(t)$ do not reach local minima for velocity and flow, nor do they reach local maxima for density. Instead, the responses exhibit local maxima for velocity and flow, and local minima for density. This is due to the presence of positive increments $\Delta x_{i}(t,\tau)$ in Eq. (5). It represents vehicles accelerating during the congestion phase, which increases flow and reduces local density. Consequently, the correlations between the three observables remain, but appear in an opposite manner. However, the results depicted in Fig. 4 indicate a difference between the phases of congestion occurring at interchange Breitscheid and those occurring at interchange Herne and on motorway A3. Besides the acceleration phase captured by $\varepsilon^{(2)}(t)$, we find a similar case in the flow response for $\varepsilon^{(0)}(t)$ at interchange Breitscheid in Fig. 4(c). A few minutes after the minimum, the response to $\varepsilon^{(0)}(t)$ propagates into a local maximum, indicating that vehicles are accelerating out of the traffic jam. Unfortunately, this only manifests in the flow response but not in the velocity or density responses to $\varepsilon^{(0)}(t)$. Although there are no distinct local maxima, there is an increasing trend in both responses. Overall, response functions are capable of capturing acceleration phases during congestion, which is an important finding.